A cable 22 feet long runs from the top of a utility pole to a point on the ground 16 feet from the base of the pole. How tall is the utility pole?
Answer by nyc_function(2741)

Let h = height of utility pole
h^2 + b^2 = c^2
Let b = 16
Lett c = 22
h^2 + (16)^2 = (22)^2
Can you take it from here?
HINT: Solve for h